. Even if glass is oven-dried, the SiO2 still has surface OH groups. Is this a concern in our Grignard reaction? Given that the surface density of hydroxyl groups of glass is about 4 OH groups/nm2, how much would your yield of Grignard reagent be reduced by surface OH groups, presuming complete reaction with these groups? Show your calculation. Assume that the glass surface area in contact with your solution is 5 cm2.
Answer #1

The use of other solid materials as catalyst supports is also common in Ziegler-Natta catalysis. Generally these materials in5 cm2 has 5e+14nm2 . So 5 cm2 will have 4 x 5e+14 = 20e+14 OH- groups which will react with 1:1 ratio with grignard reagent. So number of grignard molecule 20e+14 .

Moles\ of\ grignard =\frac{20\times 10^{14}}{6.022\times 10^{23}}=3.21\times 10^{-9}